OpenStudy (caroline1218):

Which function below represents the arithmetic sequence 3, 7, 11, 15...?

OpenStudy (caroline1218):

A. f(n)=4+3(n-1) B. f(n)=4+3n C. f(n)=3+4n D. f(n)=3+4(n-1)

OpenStudy (anonymous):

are you with me

OpenStudy (anonymous):

okay lets do this together

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

you are not to give direct answers

OpenStudy (anonymous):

Try n = 1 and n = 2 using D.

OpenStudy (caroline1218):

Ok

OpenStudy (anonymous):

exactly that is how you teach someone @robtobey

OpenStudy (anonymous):

but don't give a direct answer